# Comparison of Nested Geometry Treatments within GPU-Based Monte Carlo Neutron Transport Simulations of Fission Reactors

### v1

@@ Abstract
Monte Carlo (MC) neutron transport provides detailed estimates of radiological quantities within fission reactors. This method involves tracking individual neutrons through a computational geometry. CPU-based MC codes use multiple polymorphic tracker types with different tracking algorithms to exploit the repeated configurations of reactors, but virtual function calls have high overhead on the GPU. The Shift MC code was modified to support GPU-based tracking with three strategies: (1) dynamic polymorphism (DP) with virtual functions, (2) static polymorphism (SP), and (3) a single tracker (ST) type with tree-based acceleration. Results on the Frontier supercomputer show that the DP, SP, and ST methods achieve 77.8%, 91.2%, and 83.4% of the practical maximum tracking rate in the worst case, indicating that any of these methods can be used without incurring a significant performance penalty. The flexibility of the ST method is highlighted with a hexagonal-grid microreactor problem, performed without hexagonal-grid-specific tracking routines.
@@ 1 Introduction
Nuclear reactors account for nearly 20% of electricity production in the United States, with lifecycle greenhouse gas emissions 17–29$\times$ less than coal-fired power plants per unit of energy generated (Center for Sustainable Systems, 2022). These reactors derive energy from induced nuclear fission, a process in which a free neutron is captured by a heavy fuel nucleus, causing it to split into multiple lighter nuclei. Each fission releases heat, as well as additional free neutrons that may induce subsequent fissions, perpetuating the process. The pressurized water reactor (PWR)—the most common reactor design worldwide—consists of a core comprised of assemblies, each of which is comprised of fuel rods, as shown in Figure 1. Water is pumped through the core to extract heat, and electricity is generated with a thermodynamic power cycle via the expansion of steam through a turbine.
@@
Significant computational modeling and simulation is required for reactor design, licensing, and operation. Monte Carlo (MC) neutron transport is the preeminent method for obtaining high-fidelity estimates of radiological quantities because of its continuous (i.e., non-discrete) treatment of space, direction, and energy dimensions. This stochastic method involves simulating neutron histories—the circuitous paths that individual neutrons take within a reactor—using a random walk technique. By simulating a large number of histories, accurate statistical estimates of radiological quantities can be deduced. A key radiological quantity is the effective neutron multiplication factor ($k_{\mathrm{eff}}$), defined as the average number of neutrons born from fission that induce subsequent fissions. A $k_{\mathrm{eff}}$ of 1 indicates that the reactor is critical, i.e. operating at steady state, whereas $k_{\mathrm{eff}}<1$ and $k_{\mathrm{eff}}>1$, indicate that the fission rate will decrease or increase over time, respectively. Estimates of $k_{\mathrm{eff}}$ are obtained using MC via a power iteration scheme (Lieberoth, 1968).
@@
Figure 2: Demonstration of the CSG construction process, shown as a 2D slice. Plane surfaces (orange), have surface normals (yellow), denoting which side of the plane is positive. The outside of the blue cylindrical surface is considered positive. Cell construction logic is used to form cells 1 and 2. The cell 2 logic can be read as “the intersection of the space on the positive side of surface 1, the negative side of surface 2, the positive side of surface 3, … etc.”
@@
The nested and repeated structures that comprise reactors receive special treatment for performance and user convenience. A CSG universe is a contiguous geometric region comprised of one or more cells that can be embedded within one or more parent cells (West et al., 1979). Array universes are special universe types in which cells comprise a structured mesh. The cells in an array universe must be filled with other universes which may be either CSG universes or other array universes. Examples of CSG universes embedded in rectilinear and hexagonal array universes are shown in Figure 3. Some fission reactors, including PWRs, can be modeled with three levels of universes: an array universe representing the core, with each core array cell filled with an assembly array universe, and each assembly array cell filled with a CSG universe containing fuel rods surrounded by coolant (referred to as a pin within this work).
@@
Figure 3: Demonstration of CSG universes nested within rectilinear and hexagonal arrays universes, shown as a 2D slice.
@@
The use of array universes in concert with CSG universes provides tracking performance benefits. When determining what cell contains a given point, the hierarchical configuration of nested universes can be exploited by first finding which assembly contains the point, then which pin within the assembly, and finally which cell within the pin. This is loosely analogous to standard ray tracing acceleration structures such as bounding volume hierarchy (BVH) (Ericson, 2004) or $k$-d trees (Bentley, 1975). In addition, because CSG and array universes represent geometry differently, the tracking operations summarized in Table 1 are implemented separately for each universe type, allowing each to be optimized. Within this work, objects that carry out these tracking operations are referred to as trackers. Thus, a CSG universe has a corresponding CSG tracker, a rectilinear array universe has a rectilinear array tracker, etc. Array trackers can implement tracking operations significantly more efficiently than CSG trackers by exploiting the regular structure of array universes.
@@
Using multiple tracker types requires tracking functions to be called polymorphically. On the CPU with C++, this is accomplished by implementing tracker functions as virtual functions of a tracker base class. However, using GPU programming models (e.g., HIP, CUDA), it is not clear whether the performance benefits of different tracker types outweigh the overhead associated with virtual function calls (Zhang et al., 2021).
@@
Shift is a general purpose MC radiation transport code developed at Oak Ridge National Laboratory (ORNL) capable of simulating the behavior of neutrons and high-energy photons for fission, fusion, and national security applications (Pandya et al., 2016). Shift, written in C++ with abstracted HIP/CUDA device programming models, supports both CPU and GPU execution (Hamilton and Evans, 2019) and is designed to scale effectively from laptops to leadership-class supercomputers. On the CPU, Shift uses the Oak Ridge Adaptable Nested Geometry Engine (ORANGE) (Johnson et al., 2023) [^1] The ORANGE package is shared between Shift and the Celeritas MC code, used for high-energy physics detector analysis (Johnson et al., 2021)., which includes CSG, rectilinear array, and hexagonal array trackers, leveraged via virtual functions. On the GPU, Shift—as well other MC codes that support GPU execution such as OpenMC (Romano et al., 2015; Tramm et al., 2022), and PRAGMA (Choi et al., 2019)—have avoided GPU polymorphism by supporting only a single universe/tracker type.
@@
In this work, three methods for GPU-based multi-universe tracking were implemented in ORANGE to assess the trade-offs between single and multiple tracker types. The timing results for Shift $k_{\mathrm{eff}}$ calculations on the Summit (Oak Ridge Leadership Computing Facility, 2023b) and Frontier (Oak Ridge Leadership Computing Facility, 2023a) supercomputers at ORNL were recorded for each of the three methods and compared to RTK timing results. The three methods are outlined as follows. The dynamic polymorphism (DP) method employs multiples tracker types with virtual functions. The static polymorphism (SP) method employs multiple tracker types with switch statements. The single tracker (ST) method avoids polymorphism by using a single tracker type for all universe types. The ST method is accomplished using pseudo-array universes, which are arrays modeled as CSG universes and tracked upon with a CSG tracker. Although this approach forfeits the aforementioned benefits of array trackers, this is counteracted with a bounding interval hierarchy (BIH) acceleration structure (Wächter and Keller, 2006).
The remainder of this work proceeds as follows. Section 2 provides background on how Shift solves for $k_{\mathrm{eff}}$, including the role of tracking operations. Section 3 introduces tracking algorithms for CSG and array universes, multi-universe geometries, and RTK universes. Section 4 describes the implementation of each of the three experimental multi-universe GPU tracking methods. Section 5 describes the computer hardware on Summit and Frontier used for this work. Section 6 provides a performance comparison of the three methods using a model of the NuScale small module reactor (SMR) (NuScale Power, 2023), comprised of rectilinear assemblies. Section 7 demonstrates the flexibility of the ST method on the Empire microreactor benchmark problem (Lee et al., 2020; Matthews et al., 2021) comprised of hexagonal assemblies.
@@ 2.1 Neutron transport equation
where $\hat{\mathrm{T}}$ is the transport operator, $\hat{\mathrm{S}}$ is the scattering operator, $\hat{\mathrm{\chi}}$ is energy spectrum of neutrons born from fission, $\hat{\mathrm{F}}$ is the fission operator, and $\psi$ is the neutron flux, which describes the number of neutrons that pass through a 2D area per unit time per unit angle. The $\hat{\mathrm{T}}$, $\hat{\mathrm{S}}$, and $\hat{\mathrm{F}}$ operators depend on nuclear cross sections, but these relationships are omitted for brevity. By collecting the operators in a single term,
@@ 2.2 Monte Carlo power iteration
[…]
@@ 6         $k_{\mathrm{eff}}$, f $=$ run_histories(f, $n$)
10              sample ##\## of mean free paths ($\tau$) before event
@@ 3.1 Single universe tracking
Algorithm 3 CSG tracker version of find_cell.
@@
Algorithm 4 CSG tracker version of cross_surface.
1procedure cross_surface(cell, surf, pos)
@@
Possible implementations of find_cell and cross_surface for a rectilinear array tracker are shown Algorithms 5 and 6. These algorithms have favorable time complexity relative to their CSG counterparts. Assuming the rectilinear grid is $N^{1/3}\times N^{1/3}\times N^{1/3}$, find_cell can be performed in $O(\log(N^{1/3}))$ time by conducting a binary search over the mesh divisions in each dimension. The cross_surface operation can be performed in $O(1)$ time because the surface of each cell is known to have only one neighbor. It requires the current cell, the $i/j/k$ dimension of the surface being crossed, and whether the surface is being crossed in the positive or negative direction.
Algorithm 5 Rectilinear array tracker version of find_cell.
@@
Algorithm 6 Rectilinear array tracker version of cross_surface.
@@ 3.2 Multi-universe CPU tracking in ORANGE
The algorithms shown in Section 3.1 show how CSG and array trackers operate within single universes. Here, tracking algorithms for geometries consisting of multiple nested universes are shown, as implemented on the CPU in ORANGE. The ORANGE CPU version of find_cell and cross_surface are shown in Algorithms 7 and 8. The find_cell algorithm recursively finds the cell within daughter universes until reaching the bottom-most (i.e., most embedded) level. In cross_surface, a surf_universe argument is supplied that denotes the top-most (i.e., least embedded) level for which the neutron is on a surface, noting that it may be on a surface on multiple levels. The cell on the other side of the surface is then found by recursing through daughters starting at this level. Both algorithms assume the existence of a polymorphic get_tracker function which returns a tracker object for a given universe, depending on its type.
@@
2     tracker $=$ get_tracker(root_universe)
3     cell $=$ tracker.find_cell(pos)
@@
5         tracker $=$ get_tracker(cell.daughter)
6         cell $=$ tracker.find_cell(pos)
@@
2     tracker $=$ get_tracker(surf_universe)
3     next_cell $=$ tracker.cross_surface(args)
@@
5         tracker $=$ get_tracker(daughter)
6         next_cell $=$ tracker.find_cell(args.pos)
7         daughter $=$ cell.daughter
@@ 3.3 RTK tracking
As mentioned in Section 1, RTK is the reactor-specific universe type currently employed by Shift for GPU execution. RTK is a template universe type explicitly instantiated to contain three nested levels of universes necessary to model reactors such as PWRs with rectilinear configurations. An RTK universe consists of a rectilinear array core universe populated with rectilinear array assembly universes, each populated with pin universes. The pin universe is not a full, general purpose CSG universe, but rather a limited CSG universe consisting of concentric cylinders within a rectangular cuboid cell.
The RTK version of find_cell and cross_surface are shown in Algorithms 9 and 10. These algorithms resemble the standard multi-universe tracking algorithms shown in in Section 3.2, with two key differences. First, because there are exactly three levels of universes, the while loops over daughter universes can be unrolled (in practice, this is achieved through C++ template recursion). Second, the types of universes at each level are known at compile time. Whereas the multi-universe tracking algorithms shown in Section 3.2 rely on a polymorphic get_tracker function, Algorithms 9 and 10 can call non-polymorphic get_rect_tracker and get_csg_tracker functions. As a result of these simplifications, RTK is expected to be the most performant tracker type. However, the clear limitations on geometric complexity imposed by RTK motivate the implementation of a general-purpose GPU tracker type.
@@
2     core_tracker $=$ get_rect_tracker(core_universe)
3     core_cell $=$ core_tracker.find_cell(pos)
@@
5     assm_tracker $=$ get_rect_tracker(assm_universe)
6     assm_cell $=$ assm_tracker.find_cell(pos)
@@
8     pin_tracker $=$ get_csg_tracker(pin_universe)
9     cell $=$ pin_tracker.find_cell(pos)
@@
3         core_tracker $=$ get_rect_tracker(surf_universe)
4         core_cell $=$ core_tracker.cross_surface(args)
5         assm_universe $=$ core_cell.daughter
@@
7                $=$ get_rect_tracker(assm_universe)
8         assm_cell $=$ assm_tracker.find_cell(args.pos)
@@ 7                $=$ get_rect_tracker(assm_universe)
10         pin_tracker $=$ get_csg_tracker(pin_universe)
11         new_cell $=$ pin_tracker.find_cell(args.pos)
@@
13         assm_tracker $=$ get_arr_tracker(surf_universe)
14         assm_cell $=$ assm_tracker.find_cell(args.pos)
@@
16         pin_tracker $=$ get_csg_tracker(pin_universe)
17         new_cell $=$ pin_tracker.find_cell(args.pos)
@@
19         pin_tracker $=$ get_csg_tracker(surf_universe)
20         new_cell $=$ pin_tracker.find_cell(args.pos)

### v2

@@ Abstract
[…]
keywords:
Monte Carlo, radiation transport, GPU computing, nuclear reactor analysis, computational geometry
@@ 1 Introduction
[…]
@@
Significant computational modeling and simulation is required for reactor design, licensing, and operation. Monte Carlo (MC) neutron transport is the preeminent method for obtaining high-fidelity estimates of radiological quantities because of its continuous (i.e., non-discrete) treatment of space, direction, and energy dimensions. This stochastic method involves simulating neutron histories—the circuitous paths that individual neutrons take within a reactor—using a random walk technique. By simulating a large number of histories, accurate statistical estimates of radiological quantities can be deduced. A key radiological quantity is the effective neutron multiplication factor ($k_{\mathrm{eff}}$), defined as the average number of neutrons born from fission that induce subsequent fissions. A $k_{\mathrm{eff}}$ of 1 indicates that the reactor is critical, i.e., operating at steady state, whereas $k_{\mathrm{eff}}<1$ and $k_{\mathrm{eff}}>1$ indicate that the fission rate will decrease or increase over time, respectively. Estimates of $k_{\mathrm{eff}}$ are obtained using MC via a power iteration scheme (Lieberoth, 1968).
@@
Figure 2: Demonstration of the CSG construction process, shown as a 2D slice. Plane surfaces (orange) have surface normals (yellow) denoting which side of the plane is positive. The outside of the blue cylindrical surface is considered positive. Cell construction logic is used to form cells 1 and 2. The cell 2 logic can be read as “the intersection of the space on the positive side of surface 1, the negative side of surface 2, the positive side of surface 3, … etc.”
@@
The nested and repeated structures that comprise reactors receive special treatment for performance and user convenience. A CSG universe is a contiguous geometric region composed of one or more cells that can be embedded within one or more parent cells (West et al., 1979). Array universes are special universe types in which cells comprise a structured mesh. The cells in an array universe must be filled with other universes which may be either CSG universes or other array universes. Examples of CSG universes embedded in rectilinear and hexagonal array universes are shown in Figure 3. Some fission reactors, including PWRs, can be modeled with three levels of universes: an array universe representing the core, with each core array cell filled with an assembly array universe, and each assembly array cell filled with a CSG universe containing fuel rods surrounded by coolant (referred to as a pin within this work).
@@
Figure 3: Demonstration of CSG universes nested within rectilinear and hexagonal array universes, shown as a 2D slice. The rectilinear array universe is shown with uniform spacing for simplicity, but non-uniform spacing is required to support typical reactor applications.
@@
The use of array universes in concert with CSG universes provides tracking performance benefits. When determining what cell contains a given point, the hierarchical configuration of nested universes can be exploited by first finding which assembly contains the point, then which pin within the assembly, and finally which cell within the pin. This is loosely analogous to standard ray tracing acceleration structures such as bounding volume hierarchies (BVHs) (Ericson, 2004) or $k$-d trees (Bentley, 1975). In addition, because CSG and array universes represent geometry differently, the tracking operations summarized in Table 1 are implemented separately for each universe type, allowing each to be optimized. Within this work, objects that carry out these tracking operations are referred to as trackers. Thus, a CSG universe has a corresponding CSG tracker, a rectilinear array universe has a rectilinear array tracker, etc. Array trackers can implement tracking operations significantly more efficiently than CSG trackers by exploiting the regular structure of array universes.
@@
When using multiple tracker types, polymorphism can be employed to call tracking functions. On the CPU with C++, this is accomplished by implementing tracker functions as virtual functions of a tracker base class. However, using GPU programming models (e.g., HIP, CUDA), it is not clear whether the performance benefits of different tracker types outweigh the overhead associated with virtual function calls (Zhang et al., 2021).
@@
Shift is a general-purpose MC radiation transport code developed at Oak Ridge National Laboratory (ORNL) capable of simulating the behavior of neutrons and high-energy photons for fission, fusion, and national security applications (Pandya et al., 2016). Shift, written in C++ with abstracted HIP/CUDA device programming models, supports both CPU and GPU execution (Hamilton and Evans, 2019) and is designed to scale effectively from laptops to leadership-class supercomputers. On the CPU, Shift uses the Oak Ridge Adaptable Nested Geometry Engine (ORANGE) (Johnson et al., 2023), [^2] The ORANGE package is shared between Shift and the Celeritas MC code, used for high-energy physics detector analysis (Johnson et al., 2021). which includes CSG, rectilinear array, and hexagonal array trackers, leveraged via virtual functions. In Shift, GPU polymorphism has previously been avoided by supporting only a single universe/tracker type.
@@
In this work, three methods for GPU-based multi-universe tracking were implemented in ORANGE to assess the trade-offs between single and multiple tracker types. The timing results for Shift $k_{\mathrm{eff}}$ calculations on the Summit (Oak Ridge Leadership Computing Facility, 2023b) and Frontier (Oak Ridge Leadership Computing Facility, 2023a) supercomputers at ORNL were recorded for each of the three methods and compared to RTK timing results. The three methods are outlined as follows. The dynamic polymorphism (DP) method employs multiple tracker types with virtual functions. The static polymorphism (SP) method employs multiple tracker types with switch statements. The single tracker (ST) method avoids polymorphism by using a single tracker type for all universe types. The ST method is accomplished using pseudo-array universes (a term specific to this work), which are arrays modeled as CSG universes and tracked upon with a CSG tracker. Although this approach forfeits the aforementioned benefits of array trackers, this is counteracted with a bounding interval hierarchy (BIH) acceleration structure (Wächter and Keller, 2006).
The remainder of this work proceeds as follows. Section 2 provides background on how Shift solves for $k_{\mathrm{eff}}$, including the role of tracking operations. Section 3 introduces tracking algorithms for CSG and array universes, multi-universe geometries, and RTK universes. Section 4 describes the implementation of each of the three experimental multi-universe GPU tracking methods. Section 5 describes the computer hardware on Summit and Frontier used for this work. Section 6 provides a performance comparison of the three methods using a model of the NuScale small modular reactor (SMR) (NuScale Power, 2023), composed of rectilinear assemblies. Section 7 demonstrates the flexibility of the ST method on the Empire microreactor benchmark problem (Lee et al., 2020; Matthews et al., 2021) composed of hexagonal assemblies.
@@ 2.1 Neutron transport equation
where $\hat{\mathrm{T}}$ is the transport operator, $\hat{\mathrm{S}}$ is the scattering operator, $\hat{\mathrm{\chi}}$ is the energy spectrum of neutrons born from fission, $\hat{\mathrm{F}}$ is the fission operator, and $\psi$ is the neutron flux, which describes the number of neutrons that pass through a 2D area per unit time per unit angle. The $\hat{\mathrm{T}}$, $\hat{\mathrm{S}}$, and $\hat{\mathrm{F}}$ operators depend on nuclear cross sections, but these relationships are omitted for brevity. By collecting the operators in a single term,
@@ 2.2 Monte Carlo power iteration
where $\langle\cdot\rangle$ denotes integration over space, energy, and angle. This iteration scheme is carried out using Algorithm 1, where run_histories is an MC neutron transport simulation. As seen in Algorithm 1, iterations of Equations 5 and 6, referred to as cycles, are carried out in two stages. First, a sufficient number of inactive cycles are run to converge the shape of f, with intermediate estimates of $k_{\mathrm{eff}}$ discarded. Then active cycles are run, accumulating $k_{\mathrm{eff}}$ estimates to produce the final $k_{\mathrm{eff}}$ with statistical uncertainty. Statistical estimates of f and other radiological quantities such as neutron flux or specific reaction rates are also obtained from active cycles (omitted from Algorithms 1 and 2 for brevity). These estimates, often calculated within geometry cells or the volume elements of a superimposed mesh, are called tallies. Active cycles typically require longer compute times than inactive cycles because of the additional overhead associated with tallies.
@@ 6   $k_{\mathrm{eff}}$, f == run_histories(f, $n$)
10     sample #\# of mean free paths ($\tau$) before event
@@ 3.1 Single universe tracking
Algorithm 3 Possible CSG tracker version of find_cell.
@@
Algorithm 4 Possible CSG tracker version of cross_surface.
1procedure cross_surface(pos, cell, surf)
@@
Possible implementations of find_cell and cross_surface for a rectilinear array tracker are shown in Algorithms 5 and 6. These algorithms have favorable time complexity relative to their CSG counterparts. By conducting a binary search over the mesh divisions, find_cell can be performed in $O(\log N)$ time.  [^3] For rectilinear grids with uniform spacing, find_cell can be performed in $O(1)$ time by calculating the array indices directly in each dimension. However, the presence of gaps between adjacent assemblies in most reactors necessitates the use of non-uniform spacing and therefore a full binary search. The cross_surface operation is performed in $O(1)$ time because the surface of each cell is known to have only one neighbor. It requires the current cell, the $i/j/k$ dimension of the surface being crossed, and whether the surface is being crossed in the positive or negative direction.
Algorithm 5 Possible rectilinear array tracker version of find_cell.
@@
Algorithm 6 Possible rectilinear array tracker version of cross_surface.
@@ 3.2 Multi-universe CPU tracking in ORANGE
The algorithms shown in Section 3.1 show how possible implementations of CSG and array trackers operate within single universes. Here, tracking algorithms for geometries consisting of multiple nested universes are shown, as implemented on the CPU in ORANGE. The ORANGE CPU version of find_cell and cross_surface are shown in Algorithms 7 and 8. The find_cell algorithm recursively finds the cell within daughter universes until reaching the bottom-most (i.e., most embedded) level. In cross_surface, a surf_universe argument is supplied that denotes the top-most (i.e., least embedded) level for which the neutron is on a surface, noting that lower-level coincident boundary surfaces are removed during preprocessing. The cell on the other side of the surface is then found by recursing through daughters starting at this level. Both algorithms assume the existence of a polymorphic get_tracker function which returns a tracker object for a given universe, depending on its type.
@@
2  tracker == get_tracker(root_universe)
3  cell == tracker.find_cell(pos)
@@
5   tracker == get_tracker(cell.daughter)
6   cell == tracker.find_cell(pos)
@@
2  tracker == get_tracker(surf_universe)
3  next_cell == tracker.cross_surface(args)
@@
5   tracker == get_tracker(next_cell.daughter)
6   next_cell == tracker.find_cell(args.pos)
@@ 3.3 RTK tracking
As mentioned in Section 1, RTK is the reactor-specific universe type currently employed by Shift for GPU execution. RTK is a template universe type explicitly instantiated to contain the three nested levels of universes necessary to model reactors such as PWRs with rectilinear configurations. An RTK universe consists of a rectilinear array core universe populated with rectilinear array assembly universes, each populated with pin universes. The pin universe is not a full, general-purpose CSG universe, but rather a limited CSG universe consisting of concentric cylinders within a rectangular cuboid cell.
The RTK version of find_cell and cross_surface are shown in Algorithms 9 and 10. These algorithms resemble the standard multi-universe tracking algorithms shown in Section 3.2, with two key differences. First, because there are exactly three levels of universes, the while loops over daughter universes can be unrolled (in practice, this is achieved through C++ template recursion). Second, the types of universes at each level are known at compile time. Whereas the multi-universe tracking algorithms shown in Section 3.2 rely on a polymorphic get_tracker function, Algorithms 9 and 10 can call non-polymorphic get_rect_tracker and get_csg_tracker functions. As a result of these simplifications, RTK is expected to be the most performant tracker type. However, the clear limitations on geometric complexity imposed by RTK motivate the implementation of general-purpose GPU tracking capabilities.
@@
2  core_tracker == get_rect_tracker(core_universe)
3  core_cell == core_tracker.find_cell(pos)
@@
5  assm_tracker == get_rect_tracker(assm_universe)
6  assm_cell == assm_tracker.find_cell(pos)
@@
8  pin_tracker == get_csg_tracker(pin_universe)
9  cell == pin_tracker.find_cell(pos)
@@
3   core_tracker == get_rect_tracker(surf_universe)
4   core_cell == core_tracker.cross_surface(args)
5   assm_universe == core_cell.daughter
@@
7      == get_rect_tracker(assm_universe)
8   assm_cell == assm_tracker.find_cell(args.pos)
@@
10   pin_tracker == get_csg_tracker(pin_universe)
11   new_cell == pin_tracker.find_cell(args.pos)
@@
13   assm_tracker == get_arr_tracker(surf_universe)
14   assm_cell == assm_tracker.find_cell(args.pos)
@@
16   pin_tracker == get_csg_tracker(pin_universe)
17   new_cell == pin_tracker.find_cell(args.pos)
@@
19   pin_tracker == get_csg_tracker(surf_universe)
20   new_cell == pin_tracker.find_cell(args.pos)